OpenStudy (anonymous):

calculus help.... Which of the following families of functions is the solution to the differential equation? y' = 10xy

OpenStudy (anonymous):

y = Cx^1/10......y=x^10+C......Ce^(5x^2)

OpenStudy (anonymous):

i believe you can do separation of parameters so dy/dx =10xy dy/y=10xdx dy/y = ln y 10xdx = 5x^2 raise each side to e to get y = e^(5x^2) and then the constant

OpenStudy (anonymous):

because the 10xdx actually would be 5x^2 +c and raised to e is: e^(5x^2+c) = e^(5x^2)*e^(c) which e^c is just a number so you can turn that into c or whatever letter you want for a constant

OpenStudy (anonymous):

oh and separation of variable sorry, not parameters.

OpenStudy (anonymous):

that's right htanks!